# Retrieval-Augmented Generation for AI-Generated Content: A Survey

### v1

@@ Abstract
The development of Artificial Intelligence Generated Content (AIGC) has been facilitated by advancements in model algorithms, scalable foundation model architectures, and the availability of ample high-quality datasets. While AIGC has achieved remarkable performance, it still faces challenges, such as the difficulty of maintaining up-to-date and long-tail knowledge, the risk of data leakage, and the high costs associated with training and inference. Retrieval-Augmented Generation (RAG) has recently emerged as a paradigm to address such challenges. In particular, RAG introduces the information retrieval process, which enhances AIGC results by retrieving relevant objects from available data stores, leading to greater accuracy and robustness. In this paper, we comprehensively review existing efforts that integrate RAG technique into AIGC scenarios. We first classify RAG foundations according to how the retriever augments the generator. We distill the fundamental abstractions of the augmentation methodologies for various retrievers and generators. This unified perspective encompasses all RAG scenarios, illuminating advancements and pivotal technologies that help with potential future progress. We also summarize additional enhancements methods for RAG, facilitating effective engineering and implementation of RAG systems. Then from another view, we survey on practical applications of RAG across different modalities and tasks, offering valuable references for researchers and practitioners. Furthermore, we introduce the benchmarks for RAG, discuss the limitations of current RAG systems, and suggest potential directions for future research. Project Repo: https://github.com/hymie122/RAG-Survey.
@@ I-A Background
Recent years have witnessed a surge in interest surrounding Artificial Intelligence Generated Content (AIGC). Various content generation tools have been meticulously crafted to produce diverse objects across various modalities, such as Large Language Models (LLM) including the GPT series [1, 2, 3] and the LLAMA series [4, 5, 6] for texts and codes, DALL-E [7, 8, 9] and Stable Diffusion [10] for images, and Sora [11] for videos. The word “AIGC” emphasizes that the contents are produced by advanced generative models other than human beings or rule-based approaches. These generative models have achieved remarkable performance due to their utilization of novel model algorithms, expansive foundation model architectures, and massive high-quality datasets. Specifically, sequence-to-sequence tasks have transitioned from utilizing Long Short-Term Memory (LSTM) [12] networks to Transformer-based models [13], while image-generation tasks have shifted from Generative Adversarial Networks (GAN) [14] to latent diffusion models [10]. Notably, the architecture of foundation models, initially constituted by millions of parameters [15, 16], has now grown to incorporate billions of parameters [4, 1]. These advancements are further bolstered by the availability of rich, high-quality datasets [1, 17], which provide ample training examples to fully optimize model parameters.
Information retrieval is another pivotal application within the field of computer science. Different from generation, retrieval aims to locate relevant existing objects from a vast pool of resources. The most prevalent application of retrieval lies in web search engines, which primarily focus on the task of document retrieval [18, 19]. In the present era, efficient information retrieval systems can handle document collections on the order of billions [20, 21]. Besides documents, retrieval has also been applied for other modalities [22, 23, 24, 25].
@@
A typical RAG process is shown in Fig. 1: given an input query, the retriever locates and looks up relevant data sources, then the retrieved results interact with the generator to enhance the overall generation process. The retrieval results can interact with the generation process in different ways: they can serve as augmented input to the generator [31, 32]; they can join at the middle stage of generation as latent representations [33, 34]; they can contribute to the final generation results in the form of logits [35, 36]; they can even influence or omit certain generation steps [30, 37]. Moreover, built upon the typical, foundational RAG process, numerous enhancements have also been proposed to improve the overall quality. These enhancements encompass methods for specific components as well as optimizations targeting the entire pipeline.
@@
Although the concept of RAG initially emerged in text-to-text generation [32], it has also been adapted to various domains, including codes [38, 39, 40], audios [41, 42], images [43, 44, 45], videos [46, 47], 3D [48, 49], knowledge [50, 51, 52], and AI for science [53, 54]. In particular, the essential idea and process of RAG are largely consistent across modalities. However, it necessitates specific minor adjustments in augmentation techniques, and the selection of retrievers and generators varies depending on the specific modalities and tasks.
Despite recent rapid growth in research on RAG, a systematic review encompassing all foundations, enhancements, and applications is notably absent. The absence of discussion on RAG foundations significantly undermines the practical value of research in this domain, leaving the potential of RAG not fully explored. While the majority of research interest, particularly among LLM researchers, centers on query-based RAG in text-generation tasks, it is essential to acknowledge that other RAG foundation paradigms are also effective techniques with significant potential for usage and further development. The lack of an overview on RAG applications causes researchers and practitioners to overlook RAG’s progress across multiple modalities and remain unaware of how RAG can be effectively applied. Although text generation is typically considered as the main application of RAG, we emphasize that the development of RAG in other modalities also began at an early stage and has yielded promising advancements. Certain modalities have a rich historical connection to the retrieval technique, infusing RAG with distinctive characteristics. Despite that some surveys on RAG have already been proposed, they mainly focus on partial aspects, such as specific methodologies and tasks. In this paper, our objective is to present a comprehensive survey to provide a systematic overview of RAG.
@@ II-B1 Transformer Model
$Q=x_{in}*W_{q}+b_{q},K=x_{in}*W_{k}+b_{k},V=x_{in}*W_{v}+b_{v},$
$x_{out}=LayerNorm1(Softmax(\frac{Q*K^{T}}{\sqrt{h}})*V*W_{o}+b_{o})+x_{in},$
$y=LayerNorm2((x_{out}*W_{1}+b_{1})*W_{2}+b_{2})+x_{out}.$
@@ II-B2 LSTM
Long Short-Term Memory (LSTM) [61] is a special Recurrent Neural Network (RNN) model that overcomes the exploding/vanishing gradient problems of RNN in processing long-term dependency information by introducing cell state and gate mechanisms. The LSTM model consists of four gates: Input Gate, Forget Gate, Output Gate, and a Cell State. These gates update the cell state by controlling the information flow, enabling the model to remember long-term dependent information. Cell State is the core module of the LSTM model which can memorize and maintain information. The Input Gate decides which input data should be retained in the cell state. Forget Gate determines which cell state information should be discarded to avoid excessive memory. Output Gate determines how the information in the cell state affects the current output. The flow of data and the collaborative work process between components are shown in the Fig. 3.
@@ III-A1 Query-based RAG
Recently, some methods have proposed to implement RAG without modifying the language model architecture, which is particularly suitable for scenarios where language models are accessed through APIs. REPLUG [127] illustrates this methodology by treating the language model as a ”black box,” utilizing Contriever to seamlessly incorporate relevant external documents into the query. REPLUG LSR, a variant with LM-Supervised Retrieval, further refines this process by optimizing retrieval through language model-supervised insights, aiming to reduce perplexity scores and improve model performance by enriching its contextual understanding. RALM [128] uses the BM25 algorithm for document retrieval and predictive reranking to select pertinent documents for input integration.
@@
In the scientific domain of RAG technology, GeneGPT [54] is utilized to address genomics inquiries, whereas Chat-Orthopedist [136] aims to provide support for shared decision-making among adolescents with idiopathic scoliosis. Both approaches enhance the application effectiveness and information accuracy of large language models by integrating retrieved information into the prompts of these models.
@@ III-A2 Latent Representation-based RAG
The FiD [33] technique leverages both BM25 and DPR for sourcing supportive paragraphs. It concatenates each retrieved paragraph and its title with the query, processing them individually through the encoder. FiD reduces computational complexity and efficiently utilizes relevant information to generate answers by fusing information from multiple retrieved paragraphs in the decoder, rather than processing each paragraph in the encoder. The application of Fusion-in-Decoder methodologies transcends the realm of textual content processing, demonstrating substantial potential and adaptability in processing code, structured knowledge, and diverse multimodal datasets. Specifically within the code-related domain, technologies such as EDITSUM [138], BASHEXPLAINER [139], and RetrieveNEdit [140] adopt the FiD approach, facilitating integration through encoder-processed fusion. Re2Com [141] and RACE [141] , among other methods, also feature the design of multiple encoders for different types of inputs. In the field of Knowledge Base Question Answering (KBQA), the FiD method has been widely adopted, demonstrating significant effectiveness. UniK-QA [142], DECAF [143], SKP [144], KD-CoT [145], and ReSKGC [146] have effectively enhanced the performance of QA systems through the application of Fusion-in-Decoder technology. This illustrates that by integrating RAG for KBQA, the efficiency and accuracy of QA systems can be significantly improved. In the field of Science, RetMolRetMol [53]and DeepICL [147] employ the Fusion-in-Decoder strategy, integrating information at the decoder stage to enhance the relevance and quality of the generated molecular structures.
@@
Distinctly, TOME [153] shifts the focus towards comprehensive mention encodings, prioritizing the granularity of mention over mere entity representations. It meticulously generates encodings for each entity mention across Wikipedia, populating a repository with approximately 150 million entries. This repository, encompassing key and value encodings alongside entity IDs, empowers the retrieval of much more fine-grained information. TOME integrates an initial transformer block to process input texts, followed by TOME blocks with memory attention layers, facilitating the synthesis of multifaceted information sources and enhancing inferential reasoning capabilities, even for unencountered entities. Memorizing Transformers [29] revolutionize long document processing through the integration of a kNN-augmented attention mechanism within a Transformer layer. This innovation triggers a kNN search amidst input sequence processing, fetching data based on similarities between the sequence and stored key-value pairs, thereby elevating performance without necessitating complete retraining. This approach not only bolsters processing efficiency but also broadens the model’s memory span, enabling self-retrieval from its generated outputs and fine-tuning for extensive knowledge bases or code repositories. Unlimiformer [154], by embedding a k-nearest neighbors (kNN) index within a pre-trained encoder-decoder transformer framework, pioneers handling inputs of indefinite length. Storing hidden states of input tokens in the kNN index allows for the efficient retrieval of highly relevant tokens during decoding. This innovation extends the model’s capacity to manage prolonged sequences.
Diverging from prior methods for knowledge, EaE [155] empowers language models to internalize explicit entity knowledge. EaE introduces an entity-specific parameterization, optimizing inference efficacy through an entity memory layer embedded within the transformer architecture. This layer directly acquires entity representations from textual data, utilizing a sparse retrieval strategy to fetch the nearest entities based on their embeddings, thus refining the model’s comprehension through a calculated aggregation of entity-specific information.
@@ III-A4 Speculative RAG
Speculative RAG looks for opportunities to use retrieval instead of generation to save resources and accelerate response speed. REST [30] replaces the small models in speculative decoding [163] with retrieval, so as to generate drafts. GPTCache [37] tries to solve the problem of high latency when using the API of LLMs by building a semantic cache for storing LLM responses.
@@ III-B1 Input Enhancement
Query transformation can enhance the result of retrieval by modifying the input query. Query2doc [164] and HyDE [165] first use query to generate a pseudo document, and then use this document as a key for retrieval. The advantage of doing so is that the pseudo document will contain more rich relevant information, which helps to retrieve more accurate results.
@@
Data augmentation refers to improving the data in advance before retrieval, such as removing irrelevant information, eliminating ambiguity, updating outdated documents, synthesize new data, etc. which can effectively improve the performance of the final RAG system. MakeAnAudio [42] uses captioning and audio-text retrieval to generate caption for language-free audio to mitigate data sparsity, and adds random concept audio to improve the original audio.
@@ III-B2 Retriever Enhancement
Recursive retrieve is the process of splitting a query before retrieval, and performing multiple searches to retrieve more and higher quality content. Jagerman et al. [166] uses Chain-of-Thought(COT) [167] to enable the model to break queries down gradually and provide more rich related knowledge. LLMCS [168] applies this technology to the conversational system and obtained better retrieval results by rewriting the conversation record.
@@
Chunk optimization technique refers to adjusting the size of the chunk to achieve better retrieval results. Sentence-window retrieval [169] is an efficient approach that enhances retrieval by fetching small chunks of text and returning a window of relevant sentences surrounding the retrieved segment. This method ensures that the context before and after the targeted sentence is included, providing a more comprehensive understanding of the retrieved information. Auto-merge retrieval is another advanced RAG method of LlamaIndex [169] which organizes the document in a tree-like structure, with the parent node containing the content of all children nodes. For example, articles and paragraphs, as well as paragraphs and sentences, all follow a parent-child relationship. In the retrieve process, fine-grained search for children nodes ultimately returns the parent node, effectively providing richer information.
@@
In addition, for embedding models that already have good expression power, we can still finetune them using high-quality domain data or task related data to improve their performance in specific domains or tasks. REPLUG [127] treats LM as a black box and update the retriever model based on the final results. APICoder [129] finetunes the retriever with python files and api names, signature, description. EDITSUM [138] finetunes the retriever to decrease the jaccard distance between summaries after retrieval. SYNCHROMESH [120] adds tree distance os ASTs in the loss and uses Target Similarity Tuning to finetune the Retriever. R-ConvED [46] finetunes the Retriever with the same data as generator.
@@
Hybrid retrieve refers to the simultaneous use of multiple types of retrieval methods. RAP-Gen [174] and ReACC [132] use both dense retriever and sparse retriever to improve the quality of retrieval. Rencos [119] uses sparse retriever to retrieve similar code snippets on syntactic-level and usse dense retriever to retrieve similar code snippets on semantic-level. BASHEXPLAINER [139] first uses dense retriever to capture semantic information and then uses sparse retriever to acquire lexical information. RetDream [48] first retrieves with text and then retrieves with the image embedding.
@@
The Rerank technique refers to reordering the retrieved content in order to achieve greater diversity and better results. Re2G [175] applies a re-ranker [176] model after the traditional retriever. The effect of the re-ranker model is to re-rank retrieved documents, the purpose of which is to reduce the impact of information loss caused by compressing text into vectors on the quality of retrieval. AceCoder [177] rerank the retrieved programs with a selector. The purpose of introducing the selector is to reduce redundant programs and obtain diverse retrieved programs. XRICL [178] uses a distillation-based exemplar reranker after retrieval.
Meta-data Filtering
Meta-data filtering  [179] is another method to help processing retrieved documents which uses metadata (such as time, purpose, etc.) to filter the retrieved documents for better results.

### v2

@@ Abstract
The development of Artificial Intelligence Generated Content (AIGC) has been facilitated by advancements in model algorithms, the increasing scale of foundation models, and the availability of ample high-quality datasets. While AIGC has achieved remarkable performance, it still faces several challenges, such as the difficulty of maintaining up-to-date and long-tail knowledge, the risk of data leakage, and the high costs associated with training and inference. Retrieval-Augmented Generation (RAG) has recently emerged as a paradigm to address such challenges. In particular, RAG introduces the information retrieval process, which enhances the generation process by retrieving relevant objects from available data stores, leading to higher accuracy and better robustness. In this paper, we comprehensively review existing efforts that integrate RAG technique into AIGC scenarios. We first classify RAG foundations according to how the retriever augments the generator, distilling the fundamental abstractions of the augmentation methodologies for various retrievers and generators. This unified perspective encompasses all RAG scenarios, illuminating advancements and pivotal technologies that help with potential future progress. We also summarize additional enhancements methods for RAG, facilitating effective engineering and implementation of RAG systems. Then from another view, we survey on practical applications of RAG across different modalities and tasks, offering valuable references for researchers and practitioners. Furthermore, we introduce the benchmarks for RAG, discuss the limitations of current RAG systems, and suggest potential directions for future research. Project Repo: https://github.com/hymie122/RAG-Survey.
@@ I-A Background
Recent years have witnessed the surge in interests surrounding Artificial Intelligence Generated Content (AIGC). Various content generation tools have been meticulously crafted to produce diverse outputs across various modalities, such as Large Language Models (LLMs) including the GPT series [1, 2, 3] and the LLAMA series [4, 5, 6] for texts and codes, DALL-E [7, 8, 9] and Stable Diffusion [10] for images, and Sora [11] for videos. The word “AIGC” emphasizes that the contents are produced by advanced generative models other than human beings or rule-based approaches. These generative models have achieved remarkable performance due to the utilization of novel model algorithms, explosive scale of foundation models, and massive high-quality datasets. Specifically, sequence-to-sequence tasks have transitioned from utilizing Long Short-Term Memory (LSTM) networks [12] to Transformer-based models [13], and image-generation tasks have shifted from Generative Adversarial Networks (GANs) [14] to Latent Diffusion Models (LDMs) [10] as well. Notably, the architecture of foundation models, initially constituted by millions of parameters [15, 16], has now grown to billions or even trillions of parameters [4, 1, 17]. These advancements are further bolstered by the availability of rich, high-quality datasets [1, 18], which provide ample training samples to fully optimize model parameters.
Information retrieval is another pivotal application within the field of computer science. Different from generation, retrieval aims to locate relevant existing objects from a vast pool of resources. The most prevalent application of retrieval lies in web search engines, which primarily focus on the task of document retrieval [19, 20]. In the present era, efficient information retrieval systems can handle document collections on the order of billions [21, 22]. Besides documents, retrieval has also been applied for many other modalities [23, 24, 25, 26].
@@
A typical RAG process is shown in Fig. 1. Given an input query, the retriever locates and looks up relevant data sources, then the retrieved results interact with the generator to enhance the overall generation process. The retrieval results can interact with the generation process in different ways: they can serve as augmented input to the generator [32, 33]; they can join at the middle stage of generation as latent representations [34, 35]; they can contribute to the final generation results in the form of logits [36, 37]; they can even influence or omit certain generation steps [31, 38]. Moreover, built upon the typical, foundational RAG process, numerous enhancements have also been proposed to improve the overall quality. These enhancements encompass methods for specific components as well as optimizations targeting the entire pipeline.
@@
In addition, although the concept of RAG is initially emerged in text-to-text generation [33], it has also been adapted to various domains, including codes [39, 40, 41], audios [42, 43], images [44, 45, 46], videos [47, 48], 3D [49, 50], knowledge [51, 52, 53], and AI for science [54, 55]. In particular, the essential idea and process of RAG are largely consistent across modalities. However, it necessitates specific minor adjustments in augmentation techniques, and the selection of retrievers and generators varies depending on the specific modalities and/or applications.
Despite the rapid growth in recent research on RAG, a systematic review encompassing all foundations, enhancements, and applications is notably absent. The absence of discussion on RAG foundations significantly undermines the practical value of research in this domain, leaving the potential of RAG not fully explored. While the majority of research interest, particularly among LLM researchers, centers on query-based RAG in text-generation tasks, it is essential to acknowledge that other RAG foundation paradigms are also effective techniques with significant potential for usage and further development. The lack of an overview on RAG applications causes researchers and practitioners to overlook RAG’s progress across multiple modalities and remain unaware of how RAG can be effectively applied. Although text generation is typically considered as the main application of RAG, we emphasize that the development of RAG in other modalities also began at an early stage and has yielded promising advancements. Certain modalities have a rich historical connection to the retrieval technique, infusing RAG with distinctive characteristics. Despite that some surveys on RAG have already been proposed, they mainly focus on partial aspects, such as specific methodologies and tasks. In this paper, our objective is to present a comprehensive survey to provide a systematic overview of RAG.
@@ II-B1 Transformer Model
$Q=x_{in}*W_{q}+b_{q}$
$K=x_{in}*W_{k}+b_{k}$
$V=x_{in}*W_{v}+b_{v}$
$x_{out}=LayerNorm1(Softmax(\frac{Q*K^{T}}{\sqrt{h}})*V*W_{o}+b_{o})+x_{in}$
$y=LayerNorm2((x_{out}*W_{1}+b_{1})*W_{2}+b_{2})+x_{out}$
@@ II-B2 LSTM
Long Short-Term Memory (LSTM) [62] is a special Recurrent Neural Network (RNN) model that overcomes the exploding/vanishing gradient problems of RNN in processing long-term dependency information by introducing cell state and gate mechanisms. The LSTM model consists of three gates: Input Gate, Forget Gate, and Output Gate. These gates update the cell state by controlling the information flow, enabling the model to remember long-term dependent information. Cell State is the core module of the LSTM model which can memorize and maintain information. The Input Gate decides which input data should be retained in the cell state. Forget Gate determines which cell state information should be discarded to avoid excessive memory. Output Gate determines how the information in the cell state affects the current output. The flow of data and the collaborative work process between components are shown in the Fig. 3. The entire process can be defined as follows:
$f=sigmoid(W^{f}*x^{t}+U^{f}*y^{t-1}+b^{f})$
$z=tanh(W^{z}*x^{t}+U^{z}*y^{t-1}+b^{z})$
$i=sigmoid(W^{i}*x^{t}+U^{i}*y^{t-1}+b^{i})$
$o=sigmoid(W^{o}*x^{t}+U^{o}*y^{t-1}+b^{o})$
$c^{t}=z\odot i+f\odot c^{t-1}$
$y^{t}=o\odot tanh(c^{t})$
@@ III-A1 Query-based RAG
Recently, some methods have proposed to implement RAG without modifying the language model architecture, which is particularly suitable for scenarios where language models are accessed through APIs. REPLUG [128] illustrates this methodology by treating the language model as a ”black box,” utilizing Contriever to seamlessly incorporate relevant external documents into the query. REPLUG LSR, a variant with LM-Supervised Retrieval, further refines this process by optimizing retrieval through language model-supervised insights, aiming to reduce perplexity scores and improve model performance by enriching its contextual understanding. In-Context RALM [129] uses the BM25 algorithm for document retrieval and predictive reranking to select pertinent documents for input integration.
@@
In the scientific domain of RAG technology, Chat-Orthopedist [137] aims to provide support for shared decision-making among adolescents with idiopathic scoliosis. This approach enhances the application effectiveness and information accuracy of large language models by integrating retrieved information into the prompts of the model.
@@ III-A2 Latent Representation-based RAG
The FiD [34] technique leverages both BM25 and DPR for sourcing supportive paragraphs. It concatenates each retrieved paragraph and its title with the query, processing them individually through the encoder. FiD reduces computational complexity and efficiently utilizes relevant information to generate answers by fusing information from multiple retrieved paragraphs in the decoder, rather than processing each paragraph in the encoder. The application of Fusion-in-Decoder methodologies transcends the realm of textual content processing, demonstrating substantial potential and adaptability in processing code, structured knowledge, and diverse multimodal datasets. Specifically within the code-related domain, technologies such as EDITSUM [139], BASHEXPLAINER [140], and RetrieveNEdit [141] adopt the FiD approach, facilitating integration through encoder-processed fusion. Re2Com [142], and RACE [143] , among other methods, also feature the design of multiple encoders for different types of inputs. In the field of Knowledge Base Question Answering (KBQA), the FiD method has been widely adopted, demonstrating significant effectiveness. UniK-QA [144], DECAF [145], SKP [146], KD-CoT [147], and ReSKGC [148] have effectively enhanced the performance of QA systems through the application of Fusion-in-Decoder technology. This illustrates that by integrating RAG for KBQA, the efficiency and accuracy of QA systems can be significantly improved. In the field of Science, RetMol [54] employ the Fusion-in-Decoder strategy, integrating information at the decoder stage to enhance the relevance and quality of the generated molecular structures.
@@
Diverging from prior methods for knowledge, EaE [154] empowers language models to internalize explicit entity knowledge. EaE introduces an entity-specific parameterization, optimizing inference efficacy through an entity memory layer embedded within the transformer architecture. This layer directly acquires entity representations from textual data, utilizing a sparse retrieval strategy to fetch the nearest entities based on their embeddings, thus refining the model’s comprehension through a calculated aggregation of entity-specific information. Distinctly, TOME [155] shifts the focus towards comprehensive mention encodings, prioritizing the granularity of mention over mere entity representations. It meticulously generates encodings for each entity mention across Wikipedia, populating a repository with approximately 150 million entries. This repository, encompassing key and value encodings alongside entity IDs, empowers the retrieval of much more fine-grained information. TOME integrates an initial transformer block to process input texts, followed by TOME blocks with memory attention layers, facilitating the synthesis of multifaceted information sources and enhancing inferential reasoning capabilities, even for unencountered entities. Memorizing Transformers [30] revolutionize long document processing through the integration of a kNN-augmented attention mechanism within a Transformer layer. This innovation triggers a kNN search amidst input sequence processing, fetching data based on similarities between the sequence and stored key-value pairs, thereby elevating performance without necessitating complete retraining. This approach not only bolsters processing efficiency but also broadens the model’s memory span, enabling self-retrieval from its generated outputs and fine-tuning for extensive knowledge bases or code repositories. Unlimiformer [156], by embedding a k-nearest neighbors (kNN) index within a pre-trained encoder-decoder transformer framework, pioneers handling inputs of indefinite length. Storing hidden states of input tokens in the kNN index allows for the efficient retrieval of highly relevant tokens during decoding. This innovation extends the model’s capacity to manage prolonged sequences.
@@ III-A4 Speculative RAG
Speculative RAG looks for opportunities to use retrieval instead of generation to save resources and accelerate response speed. REST [31] replaces the small models in speculative decoding [164] with retrieval, so as to generate drafts. GPTCache [38] tries to solve the problem of high latency when using the API of LLMs by building a semantic cache for storing LLM responses. COG [165] decomposes the process of text generation into a series of copy-and-paste operations, which retrieves words or phrases from the retrieved documents instaed of generation. Cao et al. [166] proposes a new paradigm to eliminate the dependence of the final result on the quantity and quality of the first stage retrieve content, which replaces generation by directly retrieving phrase level content.
@@ III-B1 Input Enhancement
Query transformation can enhance the result of retrieval by modifying the input query. Query2doc [167] and HyDE [168] first use query to generate a pseudo document, and then use this document as a key for retrieval. The advantage of doing so is that the pseudo document will contain more rich relevant information, which helps to retrieve more accurate results. TOC [169] employs recursive retrieval augmented clarification on ambiguous questions to construct a tree of disambiguated questions, thereby generating comprehensive answers to these ambiguous questions. During the construction process of this tree structure, TOC utilizes a self-verification pruning method to ensure the factual relevance of each node.
@@
Data augmentation refers to improving the data in advance before retrieval, such as removing irrelevant information, eliminating ambiguity, updating outdated documents, synthesize new data, etc. which can effectively improve the performance of the final RAG system. MakeAnAudio [43] uses captioning and audio-text retrieval to generate caption for language-free audio to mitigate data sparsity, and adds random concept audio to improve the original audio. LESS [170] strategically selects an optimal dataset for downstream tasks by analyzing gradient information, with the aim of maximizing the dataset’s impact on fine-tuning the model’s performance in response to instructional prompts.
@@ III-B2 Retriever Enhancement
Recursive retrieve is the process of splitting a query before retrieval, and performing multiple searches to retrieve more and higher quality content. Jagerman et al. [171], RAT [172] and ReACT [173] all use Chain-of-Thought(COT) [174] to enable the model to break queries down gradually and provide more rich related knowledge. LLMCS [175] applies this technology to the conversational system and obtained better retrieval results by rewriting the conversation record. ACTIVERAG [176] employs the Cognition Nexus mechanism to skillfully merge retrieved knowledge with the language model’s inherent cognitive system, thus optimizing the COT process and effectively aiding in answer generation. RATP [177] leverages Monte-Carlo Tree Search technique for multi-stage decision-making, where each node represents a thought. The system directs action choices based on scoring model feedback on these thoughts. In node expansion, it merges retrieved documents and thoughts via prompts into the language model to generate new thoughts. Kuratov et al. [178] integrates the Transformer with an RNN, utilizing the model’s intermediate outcomes as the content for retrieval. This process was executed at each layer of the Transformer, thereby significantly extending the text window’s length and effectively mitigating the issue of ”Lost in the Middle”[179].
@@
Chunk optimization technique refers to adjusting the size of the chunk to achieve better retrieval results. Sentence-window retrieval [180] is an efficient approach that enhances retrieval by fetching small chunks of text and returning a window of relevant sentences surrounding the retrieved segment. This method ensures that the context before and after the targeted sentence is included, providing a more comprehensive understanding of the retrieved information. Auto-merge retrieval is another advanced RAG method of LlamaIndex [180] which organizes the document in a tree-like structure, with the parent node containing the content of all children nodes. For example, articles and paragraphs, as well as paragraphs and sentences, all follow a parent-child relationship. In the retrieve process, fine-grained search for children nodes ultimately returns the parent node, effectively providing richer information. To address the lack of contextual information, RAPTOR [181] employs recursive embedding, clustering, and summarization of text chunks until further clustering becomes infeasible, thereby constructing a multi-level tree structure.
@@
In addition, for embedding models that already have good expression power, we can still finetune them using high-quality domain data or task related data to improve their performance in specific domains or tasks. REPLUG [128] treats LM as a black box and update the retriever model based on the final results. APICoder [130] finetunes the retriever with python files and api names, signature, description. EDITSUM [139] finetunes the retriever to decrease the jaccard distance between summaries after retrieval. SYNCHROMESH [121] adds tree distance os ASTs in the loss and uses Target Similarity Tuning to finetune the Retriever. R-ConvED [47] finetunes the Retriever with the same data as generator. Kulkarni et al. [186] applies infoNCE loss to finetune the Retriever.
@@
Hybrid retrieve denotes the concurrent employment of a diverse array of retrieval methodologies or the extraction of information from multiple distinct sources. RAP-Gen [187] and ReACC [133] use both dense retriever and sparse retriever to improve the quality of retrieval. Rencos [120] uses sparse retriever to retrieve similar code snippets on syntactic-level and usse dense retriever to retrieve similar code snippets on semantic-level. BASHEXPLAINER [140] first uses dense retriever to capture semantic information and then uses sparse retriever to acquire lexical information. RetDream [49] first retrieves with text and then retrieves with the image embedding. CRAG [188] has designed a retrieval evaluator to assess the relevance of retrieved documents to the input query, triggering three types of retrieval actions based on varying confidence levels: if deemed correct, it directly uses the retrieval results for Knowledge Refinement; if incorrect, it resorts to Web Search; and if ambiguous, it combines both approaches. To enhance performance in question-and-answer tasks, Huang et al. [189] introduces two metrics, DKS(Dense Knowledge Similarity) and RAC(Retriever as Answer Classifier), during the retrieval phase. These metrics account for both the pertinence of the answers and the applicability of the underlying knowledge. UniMS-RAG [190] introduces a novel kind of token, termed as the ’acting token’, which determines the source from which to retrieve information.
@@
The Rerank technique refers to reordering the retrieved content in order to achieve greater diversity and better results. Re2G [191] applies a re-ranker [192] model after the traditional retriever. The effect of the re-ranker model is to re-rank retrieved documents, the purpose of which is to reduce the impact of information loss caused by compressing text into vectors on the quality of retrieval. AceCoder [193] rerank the retrieved programs with a selector. The purpose of introducing the selector is to reduce redundant programs and obtain diverse retrieved programs. XRICL [194] uses a distillation-based exemplar reranker after retrieval. Rangan [195] employs the Quantized Influence Measure, assessing statistical biases between a query and a reference to evaluate the similarity of data subsets, and uses this assessment to rerank retrieval results. UDAPDR [196] utilizes large language models for the mass generation of synthetic queries at low cost, using these synthetic query sets to train domain-specific rerankers. These rerankers act as multiple teachers, and through multi-teacher knowledge distillation techniques, they create a single retriever. LLM-R [197] iteratively trains the retriever, using feedback from a frozen LLM to rank candidate documents and train the reward model. The retriever is further trained based on knowledge distillation. Each iteration of training builds upon the retriever trained in the previous cycle, facilitating iterative optimization in subsequent rounds.
Retrieval Transformation
FILCO [198] effectively filters out irrelevant content from the retrieved text chunk, leaving only the precise supporting content. This process simplifies the task for the generator, making it easier to predict the correct answer. FiD-Light [199] initially employs an encoder to convert the retrieved content into a vector, which it then compresses, resulting in a substantial reduction of latency time. RRR [200] integrates the current query with the top-k document in each round through a template, and subsequently restructures it via a pre-trained LLMs(GPT-3.5-Turbo et al.).
Others
In addition to the above optimization methods, there are also some other optimization methods for the retrieve process. For example, Meta-data filtering  [201] is a method to help processing retrieved documents which uses metadata (such as time, purpose, etc.) to filter the retrieved documents for better results. GENREAD [202] and GRG [203] introduces a novel approach where the retrieval process is supplanted or improved by prompting a LLM to generate documents in response to a given question.